OpenStudy (anonymous):

this sucks just like this place, please come on do it what is 6 over 2 minus 7

OpenStudy (anonymous):

-4...

OpenStudy (anonymous):

Will depend on how you group things, the use of English can give either of the following: \[\left( \frac{ 6 }{ 2 } \right) - 7=3-7=-4\] or \[\frac{ 6 }{ (2-7)} =\frac{ 6 }{ -5}=-\frac{6}{5}\]